# On a conjecture of Peter Neumann on fixed points in permutation groups

### v1

@@ 8.2. Class $\mathcal{C}_{2}$
The other cases are similar; we will list the choice of the elements, generally omitting the details of the calculations. If $G=\mathrm{SU}_{n}(q)$ and $M$ is of type $\mathrm{GU}_{m}(q)\wr S_{t}$, we have $(m,q)\neq(2,2)$ by [7, Proposition 2.3.6]. We choose $g=(g^{\prime},1\ldots,1)\tau$ where $g^{\prime}$ is irreducible for $m$ odd, and irreducible on a hyperplane for $m$ even, and $\tau$ is a $t$-cycle. We have that $g$ is regular and so $|C_{G}(g)|<(q+1)^{n}$ by Lemma 7.11(iv). Similarly to the $\mathrm{SL}_{n}(q)$ case we reduce to the cases $n\leq 4$ or $(n,m)=(6,3)$. If $m=n/2$ then we choose $g=g_{1}\perp g_{2}\in\mathrm{GU}_{n/2}(q)^{2}\cap G$ where $|g_{1}|=q^{3}+1$ for $n=6$ (resp. $|g_{1}|=q^{4}-1$ for $n=4$) and $g_{2}$ is diagonal; we have $\mathrm{fix}(g,G/M)=1$. The remaining cases are $(n,m)=(3,1),(4,1)$; in view of Lemma 8.1 we have $q\geq 7$ and $q\geq 5$ in the respective cases and we can choose a diagonal regular (semisimple) element $g$, such that moreover $\mathrm{fpr}(g,G/M)=1$ for $n=3$. If $G=\mathrm{Sp}_{n}(q)$ and $M$ of type $\mathrm{Sp}_{m}(q)\wr S_{t}$, then $(m,q)\neq(2,2)$ by [7, Proposition 2.3.6]. We choose $g=(g^{\prime},1,\ldots,1)\tau$ where $g^{\prime}$ is irreducible and $\tau$ is a $t$-cycle; then $g$ is regular and $|C_{G}(g)|<(q+1)^{n/2}$ (since $g$ does not have eigenvalues $\pm 1$; see the proof of Lemma 7.11(iv)). As above, we reduce to the case $(n,m)=(4,2)$. We have $q\geq 5$ by Lemma 8.1 and we can choose a regular semisimple element $g$ of $\mathrm{Sp}_{m}(q)^{t}$ of order $q+1$ with $\mathrm{fix}(g,G/M)=1$. Assume now $G=\Omega^{\varepsilon}_{n}(q)$ and $M$ of type $\mathrm{O}^{-}_{m}(q)\wr S_{t}$, so $\varepsilon=(-1)^{t}$. If $q$ is odd and $(m,q)\neq(2,3)$, we choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{\varepsilon}_{n}(q)$ where $g^{\prime}$ irreducible, so $g$ is regular. (Note $\operatorname{det}(\tau)=1$, and since $(m,q)\neq(2,3)$, we can choose $g\in\Omega^{-}_{m}(q)$ or $\mathrm{SO}^{-}_{m}(q)\smallsetminus\Omega^{-}_{m}(q)$ so that $g\in\Omega_{n}^{\varepsilon}(q)$.) If $(m,q)=(2,3)$, we write $V=W\perp V_{t}$ where $W=V_{1}\perp\cdots\perp V_{t-1}$, and accordingly we choose $g=g_{1}\perp g_{2}$ where $g_{1}=(g^{\prime},1,\ldots,1)\tau\in\mathrm{SO}^{\varepsilon}_{n-2}(q)$ is as above, and either $g_{2}=1$ or $|g_{2}|=4$, so that $g\in\Omega^{\varepsilon}_{n}(q)$. If $g_{2}=1$ then $g$ is regular; if $|g_{2}|=4$ then the dimension of the centralizer in $\mathrm{O}_{n}(K)$ is at most $n/2+2$, which gives the conclusion. Assume then $q$ is even. Then we choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{\varepsilon}_{n}(q)$ where $g^{\prime}$ is either irreducible, or irreducible on a nondegenerate codimension $2$ subspace and a reflection on the perpendicular complement (according to whether $\tau\in\Omega^{\varepsilon}_{n}(q)$ or not). Then $g$ is regular. Assume now $M$ is of type $\mathrm{O}_{m}(q)\wr S_{t}$ where $mq$ is odd and $G=\Omega^{\varepsilon}_{n}(q)$ with $\varepsilon\in\{+,-,\circ\}$. Suppose first $m\geq 3$; we have $(m,q)\neq(3,3)$ by [7, Proposition 2.3.6]. We may then choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{\varepsilon}_{n}(q)$ where $g^{\prime}$ is irreducible on a hyperplane and $\pm 1$ on the complement. Assume now $m=1$. Then the $t$-cycle $\tau$ has determinant $(-1)^{t-1}$. If $t$ is odd then choose $g=\tau\in\Omega^{\varepsilon}_{n}(q)$, so $g$ is regular. If $t$ is even then choose $g=\tau^{\prime}$ a (t1)(t-1)-cycle, so $g$ is also regular since it has at most two $1$-Jordan blocks, one of which of size one (see [56, Theorem 3.1]). Assume now $M$ is of type $\mathrm{O}_{n/2}(q)^{2}$ where $nq/2$ is odd. Then we may choose $g=g_{1}\perp g_{2}$ where $g_{1}$ and $g_{2}$ are regular and $g$ is also regular. Finally, if $M$ is of type $\mathrm{GL}_{n/2}(q^{u}).2$ with $G=\mathrm{Sp}_{n}(q),\Omega^{+}_{n}(q)$ and $u=1$, or $G=\mathrm{SU}_{n}(q)$ and $u=2$, then we may choose an irreducible element $g$ of $\mathrm{GL}_{n/2}(q^{u})$ that is regular in $G$ and the result follows. ∎
@@ 8.3. Class $\mathcal{C}_{3}$
Assume now $G=\mathrm{SU}_{n}(q)$ and $M$ of type $\mathrm{GU}_{n/r}(q^{r})$ for an odd prime divisor $r$ of $n$. If $n/r$ is odd, then let $g\in M$ be an element of order $(q^{n}+1)/(q+1)$ acting irreducibly on $V$; as above $\mathrm{fix}(g,G/M)=1$. If $n/r$ is even, we let $g\in M$ be of order $(q^{n}-1)/(q+1)$, acting irreducibly on each of a pair of complementary maximal totally singular spaces $A$ and $B$. We see that $A\not\cong B$ as $\mathbb{F}_{q^{2}}\langle g^{r}\rangle$-modules. In particular, setting $L=\mathrm{GL}_{n}(q^{2})$ we have that $C_{L}(g)=C_{L}(g^{r})=H\times K$ where $H$ and $K$ are Singer cycles on $A$ and $B$ (and $H$ and $K$ have the same eigenvalues). Then, by Lemma 8.5 the number of $r$-extension field subgroups of $L$ containing $g$ is $r$, so $\mathrm{fix}(g,G/M)\leq r$ and the result follows.
@@
Assume $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{O}^{+}_{n/r}(q^{r})$, so $n/r\geq 4$ since $M\in\mathcal{A}$. Then we let $g\in\Omega^{+}_{n/r}(q^{r})$ be of order $(q^{n/2}-1)/(2,q-1)$, acting irreducibly on each of a pair of complementary maximal totally singular subspaces $A$ and $B$. We deduce $\mathrm{fix}(g,G/M)\leq r$ as in the unitary case.
Now let $G=\Omega^{-}_{n}(q)$. If $M$ is of type $\mathrm{O}^{-}_{n/r}(q^{r})$ with $n/r\geq 4$, we let $g\in\Omega^{-}_{n/r}(q^{r})$ be of order $(q^{n/2}+1)/(2,q-1)$, and then $\mathrm{fix}(g,G/M)=1$. If $M$ is of type $\mathrm{O}_{n/r}(q^{r})$ with $n/r\geq 3$ odd, then we choose $g\in\Omega_{n/r}(q^{r})$ of order $(q^{(n-r)/2}+1)/(2,q-1)$, acting irreducibly on a nondengenerate (nr)(n-r)-space AA and trivially on $B=A^{\perp}$. Setting $L=\mathrm{GL}_{n}(q)$, we have that $C_{L}(g)=C_{L}(g^{r})=H\times\mathrm{GL}_{r}(q)$ where $H$ is a Singer cycle on $A$. Then, by Lemma 8.5 the number of $r$-extension field subgroups of $L$ containing $g$ is equal to the size of the conjugacy class of a Singer cycle in $\mathrm{GL}_{r}(q)$, which is $s:=|\mathrm{GL}_{r}(q)|/(q^{r}-1)$, so $\mathrm{fix}(g,G/M)\leq s$.
@@ 8.4. Class $\mathcal{C}_{4}$
Assume now $G=\mathrm{Sp}_{n}(q)$, so $M$ is of type $\mathrm{Sp}_{n_{1}}(q)\otimes\mathrm{O}^{\varepsilon}_{n_{2}}(q)$ with $n_{1}n_{2}=n$, $q$ odd, and $\varepsilon\in\{+,-,\circ\}$. Since $M\in\mathcal{A}$, we have $n_{2}\geq 3$ (see [39, Proposition 4.4.4]). Now we let $g=g_{1}\otimes g_{2}\in M$, where $g_{1}\in\mathrm{Sp}_{n_{1}}(q)$ is unipotent with a single Jordan block and $g_{2}\in\mathrm{O}^{\varepsilon}_{n_{2}}(q)$ is irreducible if $\varepsilon=$-, irreducible on complementary totally singular subspaces if $\varepsilon=+$, irreducible on a nondegenerate hyperplane if $\varepsilon=\circ$. Then $g$ is regular, and moreover $|C_{G}(g)|\leq 2(q+1)^{n/2}$ (since $g$ does not have both $1$ and 1-1 as eigenvalues; see the proof of Lemma 7.11(iv)). We then see that $|M||C_{G}(g)|^{3}<|G|$.
The other cases are $G=\mathrm{SU}_{n}(q)$ and $M$ of type $\mathrm{GU}_{n_{1}}(q)\otimes\mathrm{GU}_{n_{2}}(q)$; $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{Sp}_{n_{1}}(q)\otimes\mathrm{Sp}_{n_{2}}(q)$ or $\mathrm{O}^{\xi_{1}}_{n_{1}}(q)\otimes\mathrm{O}^{\xi_{2}}_{n_{2}}(q)$ (in the latter case, $q$ odd and $(\xi_{1},\xi_{2})\neq(\circ,-),(-,\circ)$); $G=\Omega_{n}(q)$ with $nq$ odd and $M$ of type $\mathrm{O}_{n_{1}}(q)\otimes\mathrm{O}_{n_{2}}(q)$; $G=\Omega^{-}_{n}(q)$ with $q$ odd and $M$ of type $\mathrm{O}_{n_{1}}(q)\otimes\mathrm{O}^{-}_{n_{2}}(q)$ with $n_{1}$ odd. Except when $M$ is of type $\mathrm{O}^{\xi_{1}}_{n_{1}}(q)\otimes\mathrm{O}^{\xi_{2}}_{n_{2}}(q)$ with $n_{1}$ and $n_{2}$ even, we can choose $g=g_{1}\otimes g_{2}$ where one element $g_{i}$ is unipotent with one Jordan block, and the other element $g_{j}$ is semisimple with distinct eigenvalues, so $g$ is regular. In the remaining case, we can choose $g_{1}$ unipotent with two Jordan blocks of size $n_{1}/2$, and $g_{2}$ semisimple with distinct eigenvalues. Then the dimension of the centralizer of $g$ in $\mathrm{O}_{n}(K)$ is $n$, and the conclusion follows. ∎
@@ 8.5. Class $\mathcal{C}_{5}$
Moreover $|G:M|=|G:M_{0}|/|M:M_{0}|$, and using $n\geq 3$ we see that $\mathrm{fix}(g,G/M)^{3}<|G:M|$. Recall now that we excluded the case $(n,q)=(4,8)$. In this case we have $M=\mathrm{SU}_{4}(2)$ and we choose $g\in M$ of order $5$. Then $g^{G}\cap M=g^{M}$ and
@@ 8.6. Class $\mathcal{C}_{6}$
Assume first $n=2$. Then, letting $E$ be the quaternion group of order $8$, we have $M=E.S_{3}=E.\mathrm{O}^{-}_{2}(2)$ or $M=E.3$. If $M=E.S_{3}$ (resp. $M=E.3$), let $g\in M$ be an element of order $3$ (resp. $4$). Then $g$ is regular in $G$ and so $|C_{G}(g)|\leq q+1$. Moreover, letting $\delta$ be the number of elements of $M$ of order $3$ (resp. $4$) in $M$, we have $\delta=8$ (resp. $\delta=6$), so $|g^{G}\cap M|\leq\delta$ and
@@
Assume next $n=3$, so $E.Q_{8}\leq M\leq E:\mathrm{Sp}_{2}(3)$, where $E$ is an extraspecial group of order $27$ and exponent $3$. We may assume $q\geq 7$ in view of Lemma 8.1. Let $g\in M$ be of order $4$, so $g$ is regular semisimple in $G$ by Lemma 8.9, and $|C_{G}(g)|\leq q^{2}-1$. The number of elements of $M$ of order $4$ is $54$. (Indeed, letting $Z$ be the center of $M$, we have $M/Z\geq\mathbb{F}_{3}^{2}\rtimes Q_{8}$, which has $9\cdot 6$ elements of order $4$, and each of these admits a unique lift in $G$ of order $4$.) Therefore
@@
since $|M|\geq 27\cdot 8$. Now assume $n=4$, so $(4\circ E).A_{6}\leq M\leq(4\circ E).\mathrm{Sp}_{4}(2)$, where $E$ is any extraspecial group of order $2^{5}$. Then by Lemma 8.9 an element of $M$ of order $5$ is regular in $G$, and a similar calculation as above suffices. Finally, assume $n\geq 5$, so $M=ZE.\mathrm{Sp}_{2m}(r)$ where $Z=Z(\mathrm{SL}_{n}(q))$ (see [39, Propositions 4.6.5 and 4.6.6]). Let $g\in M$ be of order $r^{m}+1$, so $g$ is regular in $\mathrm{GL}_{n}(K)$ by Lemma 8.9. Moreover, we have $C_{\mathrm{GL}_{n}(q)}(g)\mathrm{SL}_{n}(q)=\mathrm{GL}_{n}(q)$, and so $|C_{G}(g)|\leq(q^{n}-1)/(q-1)$. Therefore
@@
Assume now $G=\mathrm{Sp}_{n}(q)$, so $E.\Omega^{-}_{2m}(2)\leq M\leq E.\mathrm{O}^{-}_{2m}(2)$, where $n=2^{m}$ and $E=D_{8}\circ\cdots\circ D_{8}\circ Q_{8}$ is an extraspecial group of order $2^{2m+1}$. Let $g\in M$ be of order $2^{m}+1$, so $g$ is regular in $\mathrm{GL}_{n}(K)$ by Lemma 8.9, and the dimension of the centralizer of $g$ in $\mathrm{Sp}_{n}(K)$ is $n/2$, and moreover $|C_{G}(g)|<4(q+1)^{n/2}$ by Lemma 7.11(iv). For $n\geq 8$ we see that $|M||C_{G}(g)|^{3}<|G|$ and the result follows. Assume then $n=4$. We have $q\geq 5$ in view of Lemma 8.1. We have $|g|=5$, and if $p\neq 5$ then $|C_{G}(g)|\leq q^{2}+1$, and if $p=5$ then $|C_{G}(g)|=2q^{2}$. The number of elements of $E.\mathrm{O}^{-}_{4}(2)=E.S_{5}$ of order $5$ is $16\cdot 24=384$, so
@@
The other cases are similar. Assume $G=\mathrm{SU}_{n}(q)$ and $M$ of type $E.\mathrm{Sp}_{2m}(r)$ where $E$ is extraspecial of order $r^{2m+1}$ or $E=C_{4}\circ R$ where $R$ is extraspecial of order $2^{2m+1}$. Exactly as in the linear case, if $n=3$ (resp. $n=4$, resp. $n\geq 5$) then we choose $g$ of order $4$ (resp. $5$, resp. $r^{m}+1$), so $g$ is regular in $\mathrm{GL}_{n}(K)$. Finally, assume $G=\Omega^{+}_{n}(q)$ and $E.\Omega^{+}_{2m}(2)\leq M\leq E.\mathrm{O}^{+}_{2m}(2)$, where $n=2^{m}$ and $E=D_{8}\circ\cdots\circ D_{8}$ is an extraspecial group of order $2^{2m+1}$. Let $U:=E/Z(E)$ and let $g\in M$ be of order $2^{m}-1$, such that as an $\mathbb{F}_{2}\langle g\rangle$-module $U=A/Z(E)\oplus B/Z(E)$, where $A/Z(E)$ and $B/Z(E)$ are totally singular and irreducible. By Lemma 8.9, the dimension of the centralizer of $g$ in $\mathrm{GL}_{n}(K)$ is $n+2$. By Lemma 7.10(ii), it follows that the dimension of the centralizer of $g$ in $\mathrm{O}_{n}(K)$ is at most $(n+2)/2$, and we conclude with a similar calculation as above. ∎

### v2

@@ 8.2. Class $\mathcal{C}_{2}$
The other cases are similar; we will list the choice of the elements, generally omitting the details of the calculations. If $G=\mathrm{SU}_{n}(q)$ and $M$ is of type $\mathrm{GU}_{m}(q)\wr S_{t}$, we have $(m,q)\neq(2,2)$ by [7, Proposition 2.3.6]. We choose $g=(g^{\prime},1\ldots,1)\tau$ where $g^{\prime}$ is irreducible for $m$ odd, and irreducible on a hyperplane for $m$ even, and $\tau$ is a $t$-cycle. We have that $g$ is regular and so $|C_{G}(g)|<(q+1)^{n}$ by Lemma 7.11(iv). Similarly to the $\mathrm{SL}_{n}(q)$ case we reduce to the cases $n\leq 4$ or $(n,m)=(6,3)$. If $m=n/2$ then we choose $g=g_{1}\perp g_{2}\in\mathrm{GU}_{n/2}(q)^{2}\cap G$ where $|g_{1}|=q^{3}+1$ for $n=6$ (resp. $|g_{1}|=q^{4}-1$ for $n=4$) and $g_{2}$ is diagonal; we have $\mathrm{fix}(g,G/M)=1$. The remaining cases are $(n,m)=(3,1),(4,1)$; in view of Lemma 8.1 we have $q\geq 7$ and $q\geq 5$ in the respective cases and we can choose a diagonal regular (semisimple) element $g$, such that moreover $\mathrm{fpr}(g,G/M)=1$ for $n=3$. If $G=\mathrm{Sp}_{n}(q)$ and $M$ of type $\mathrm{Sp}_{m}(q)\wr S_{t}$, then $(m,q)\neq(2,2)$ by [7, Proposition 2.3.6]. We choose $g=(g^{\prime},1,\ldots,1)\tau$ where $g^{\prime}$ is irreducible and $\tau$ is a $t$-cycle; then $g$ is regular and $|C_{G}(g)|<(q+1)^{n/2}$ (since $g$ does not have eigenvalues $\pm 1$; see the proof of Lemma 7.11(iv)). As above, we reduce to the case $(n,m)=(4,2)$. We have $q\geq 5$ by Lemma 8.1 and we can choose a regular semisimple element $g$ of $\mathrm{Sp}_{m}(q)^{t}$ of order $q+1$ with $\mathrm{fix}(g,G/M)=1$. Assume now $G=\Omega^{\varepsilon}_{n}(q)$ and $M$ of type $\mathrm{O}^{-}_{m}(q)\wr S_{t}$, so $\varepsilon=(-1)^{t}$. If $q$ is odd and $(m,q)\neq(2,3)$, we choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{\varepsilon}_{n}(q)$ where $g^{\prime}$ irreducible, so $g$ is regular. (Note $\operatorname{det}(\tau)=1$, and since $(m,q)\neq(2,3)$, we can choose $g\in\Omega^{-}_{m}(q)$ or $\mathrm{SO}^{-}_{m}(q)\smallsetminus\Omega^{-}_{m}(q)$ so that $g\in\Omega_{n}^{\varepsilon}(q)$.) If $(m,q)=(2,3)$, we write $V=W\perp V_{t}$ where $W=V_{1}\perp\cdots\perp V_{t-1}$, and accordingly we choose $g=g_{1}\perp g_{2}$ where $g_{1}=(g^{\prime},1,\ldots,1)\tau\in\mathrm{SO}^{\varepsilon}_{n-2}(q)$ is as above, and either $g_{2}=1$ or $|g_{2}|=4$, so that $g\in\Omega^{\varepsilon}_{n}(q)$. If $g_{2}=1$ then $g$ is regular; if $|g_{2}|=4$ then the dimension of the centralizer in $\mathrm{O}_{n}(K)$ is at most $n/2+2$, which gives the conclusion. Assume then $q$ is even. Then we choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{\varepsilon}_{n}(q)$ where $g^{\prime}$ is either irreducible, or irreducible on a nondegenerate codimension 22 subspace and a reflection on the perpendicular complement (according to whether $\tau\in\Omega^{\varepsilon}_{n}(q)$ or not). Then $g$ is regular. Assume now $M$ is of type $\mathrm{O}_{m}(q)\wr S_{t}$ where mqmq is odd and $G=\Omega^{\varepsilon}_{n}(q)$ with $\varepsilon\in\{+,-,\circ\}$. Suppose first $m\geq 3$; we have $(m,q)\neq(3,3)$ by [7, Proposition 2.3.6]. We may then choose $g=(g^{\prime},1,\ldots,1)\tau\in\Omega^{\varepsilon}_{n}(q)$ where $g^{\prime}$ is irreducible on a hyperplane and $\pm 1$ on the complement. Assume now $m=1$. Then the $t$-cycle $\tau$ has determinant $(-1)^{t-1}$. If $t$ is odd then choose $g=\tau\in\Omega^{\varepsilon}_{n}(q)$, so $g$ is regular. If $t$ is even then choose $g=\tau^{\prime}$ a (t1)(t-1)-cycle, so $g$ is also regular since it has at most two 11-Jordan blocks, one of which of size one (see [56, Theorem 3.1]). Assume now $M$ is of type $\mathrm{O}_{n/2}(q)^{2}$ where $nq/2$ is odd. Then we may choose $g=g_{1}\perp g_{2}$ where $g_{1}$ and $g_{2}$ are regular and $g$ is also regular. Finally, if $M$ is of type $\mathrm{GL}_{n/2}(q^{u}).2$ with $G=\mathrm{Sp}_{n}(q),\Omega^{+}_{n}(q)$ and $u=1$, or $G=\mathrm{SU}_{n}(q)$ and $u=2$, then we may choose an irreducible element $g$ of $\mathrm{GL}_{n/2}(q^{u})$ that is regular in $G$ and the result follows. ∎
@@ 8.3. Class $\mathcal{C}_{3}$
Assume now $G=\mathrm{SU}_{n}(q)$ and $M$ of type $\mathrm{GU}_{n/r}(q^{r})$ for an odd prime divisor $r$ of $n$. If $n/r$ is odd, then let $g\in M$ be an element of order $(q^{n}+1)/(q+1)$ acting irreducibly on $V$; as above $\mathrm{fix}(g,G/M)=1$. If $n/r$ is even, we let $g\in M$ be of order $(q^{n}-1)/(q+1)$, acting irreducibly on each of a pair of complementary maximal totally singular spaces AA and $B$. We see that $A\not\cong B$ as $\mathbb{F}_{q^{2}}\langle g^{r}\rangle$-modules. In particular, setting $L=\mathrm{GL}_{n}(q^{2})$ we have that $C_{L}(g)=C_{L}(g^{r})=H\times K$ where $H$ and $K$ are Singer cycles on AA and $B$ (and $H$ and $K$ have the same eigenvalues). Then, by Lemma 8.5 the number of $r$-extension field subgroups of $L$ containing $g$ is $r$, so $\mathrm{fix}(g,G/M)\leq r$ and the result follows.
@@
Assume $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{O}^{+}_{n/r}(q^{r})$, so $n/r\geq 4$ since $M\in\mathcal{A}$. Then we let $g\in\Omega^{+}_{n/r}(q^{r})$ be of order $(q^{n/2}-1)/(2,q-1)$, acting irreducibly on each of a pair of complementary maximal totally singular subspaces AA and $B$. We deduce $\mathrm{fix}(g,G/M)\leq r$ as in the unitary case.
Now let $G=\Omega^{-}_{n}(q)$. If $M$ is of type $\mathrm{O}^{-}_{n/r}(q^{r})$ with $n/r\geq 4$, we let $g\in\Omega^{-}_{n/r}(q^{r})$ be of order $(q^{n/2}+1)/(2,q-1)$, and then $\mathrm{fix}(g,G/M)=1$. If $M$ is of type $\mathrm{O}_{n/r}(q^{r})$ with $n/r\geq 3$ odd, then we choose $g\in\Omega_{n/r}(q^{r})$ of order $(q^{(n-r)/2}+1)/(2,q-1)$, acting irreducibly on a nondengenerate (nr)(n-r)-space AA and trivially on $B=A^{\perp}$. Setting $L=\mathrm{GL}_{n}(q)$, we have that $C_{L}(g)=C_{L}(g^{r})=H\times\mathrm{GL}_{r}(q)$ where $H$ is a Singer cycle on AA. Then, by Lemma 8.5 the number of $r$-extension field subgroups of $L$ containing $g$ is equal to the size of the conjugacy class of a Singer cycle in $\mathrm{GL}_{r}(q)$, which is $s:=|\mathrm{GL}_{r}(q)|/(q^{r}-1)$, so $\mathrm{fix}(g,G/M)\leq s$.
@@ 8.4. Class $\mathcal{C}_{4}$
Assume now $G=\mathrm{Sp}_{n}(q)$, so $M$ is of type $\mathrm{Sp}_{n_{1}}(q)\otimes\mathrm{O}^{\varepsilon}_{n_{2}}(q)$ with $n_{1}n_{2}=n$, $q$ odd, and $\varepsilon\in\{+,-,\circ\}$. Since $M\in\mathcal{A}$, we have $n_{2}\geq 3$ (see [39, Proposition 4.4.4]). Now we let $g=g_{1}\otimes g_{2}\in M$, where $g_{1}\in\mathrm{Sp}_{n_{1}}(q)$ is unipotent with a single Jordan block and $g_{2}\in\mathrm{O}^{\varepsilon}_{n_{2}}(q)$ is irreducible if $\varepsilon=$-, irreducible on complementary totally singular subspaces if $\varepsilon=+$, irreducible on a nondegenerate hyperplane if $\varepsilon=\circ$. Then $g$ is regular, and moreover $|C_{G}(g)|\leq 2(q+1)^{n/2}$ (since $g$ does not have both 11 and 1-1 as eigenvalues; see the proof of Lemma 7.11(iv)). We then see that $|M||C_{G}(g)|^{3}<|G|$.
The other cases are $G=\mathrm{SU}_{n}(q)$ and $M$ of type $\mathrm{GU}_{n_{1}}(q)\otimes\mathrm{GU}_{n_{2}}(q)$; $G=\Omega^{+}_{n}(q)$ and $M$ of type $\mathrm{Sp}_{n_{1}}(q)\otimes\mathrm{Sp}_{n_{2}}(q)$ or $\mathrm{O}^{\xi_{1}}_{n_{1}}(q)\otimes\mathrm{O}^{\xi_{2}}_{n_{2}}(q)$ (in the latter case, $q$ odd and $(\xi_{1},\xi_{2})\neq(\circ,-),(-,\circ)$); $G=\Omega_{n}(q)$ with nqnq odd and $M$ of type $\mathrm{O}_{n_{1}}(q)\otimes\mathrm{O}_{n_{2}}(q)$; $G=\Omega^{-}_{n}(q)$ with $q$ odd and $M$ of type $\mathrm{O}_{n_{1}}(q)\otimes\mathrm{O}^{-}_{n_{2}}(q)$ with $n_{1}$ odd. Except when $M$ is of type $\mathrm{O}^{\xi_{1}}_{n_{1}}(q)\otimes\mathrm{O}^{\xi_{2}}_{n_{2}}(q)$ with $n_{1}$ and $n_{2}$ even, we can choose $g=g_{1}\otimes g_{2}$ where one element $g_{i}$ is unipotent with one Jordan block, and the other element $g_{j}$ is semisimple with distinct eigenvalues, so $g$ is regular. In the remaining case, we can choose $g_{1}$ unipotent with two Jordan blocks of size $n_{1}/2$, and $g_{2}$ semisimple with distinct eigenvalues. Then the dimension of the centralizer of $g$ in $\mathrm{O}_{n}(K)$ is $n$, and the conclusion follows. ∎
@@ 8.5. Class $\mathcal{C}_{5}$
Moreover $|G:M|=|G:M_{0}|/|M:M_{0}|$, and using $n\geq 3$ we see that $\mathrm{fix}(g,G/M)^{3}<|G:M|$. Recall now that we excluded the case $(n,q)=(4,8)$. In this case we have $M=\mathrm{SU}_{4}(2)$ and we choose $g\in M$ of order 55. Then $g^{G}\cap M=g^{M}$ and
@@ 8.6. Class $\mathcal{C}_{6}$
Assume first $n=2$. Then, letting $E$ be the quaternion group of order 88, we have $M=E.S_{3}=E.\mathrm{O}^{-}_{2}(2)$ or $M=E.3$. If $M=E.S_{3}$ (resp. $M=E.3$), let $g\in M$ be an element of order 33 (resp. 44). Then $g$ is regular in $G$ and so $|C_{G}(g)|\leq q+1$. Moreover, letting $\delta$ be the number of elements of $M$ of order 33 (resp. 44) in $M$, we have $\delta=8$ (resp. $\delta=6$), so $|g^{G}\cap M|\leq\delta$ and
@@
Assume next $n=3$, so $E.Q_{8}\leq M\leq E:\mathrm{Sp}_{2}(3)$, where $E$ is an extraspecial group of order $27$ and exponent 33. We may assume $q\geq 7$ in view of Lemma 8.1. Let $g\in M$ be of order 44, so $g$ is regular semisimple in $G$ by Lemma 8.9, and $|C_{G}(g)|\leq q^{2}-1$. The number of elements of $M$ of order 44 is $54$. (Indeed, letting $Z$ be the center of $M$, we have $M/Z\geq\mathbb{F}_{3}^{2}\rtimes Q_{8}$, which has $9\cdot 6$ elements of order 44, and each of these admits a unique lift in $G$ of order 44.) Therefore
@@
since $|M|\geq 27\cdot 8$. Now assume $n=4$, so $(4\circ E).A_{6}\leq M\leq(4\circ E).\mathrm{Sp}_{4}(2)$, where $E$ is any extraspecial group of order $2^{5}$. Then by Lemma 8.9 an element of $M$ of order 55 is regular in $G$, and a similar calculation as above suffices. Finally, assume $n\geq 5$, so $M=ZE.\mathrm{Sp}_{2m}(r)$ where $Z=Z(\mathrm{SL}_{n}(q))$ (see [39, Propositions 4.6.5 and 4.6.6]). Let $g\in M$ be of order $r^{m}+1$, so $g$ is regular in $\mathrm{GL}_{n}(K)$ by Lemma 8.9. Moreover, we have $C_{\mathrm{GL}_{n}(q)}(g)\mathrm{SL}_{n}(q)=\mathrm{GL}_{n}(q)$, and so $|C_{G}(g)|\leq(q^{n}-1)/(q-1)$. Therefore
@@
Assume now $G=\mathrm{Sp}_{n}(q)$, so $E.\Omega^{-}_{2m}(2)\leq M\leq E.\mathrm{O}^{-}_{2m}(2)$, where $n=2^{m}$ and $E=D_{8}\circ\cdots\circ D_{8}\circ Q_{8}$ is an extraspecial group of order $2^{2m+1}$. Let $g\in M$ be of order $2^{m}+1$, so $g$ is regular in $\mathrm{GL}_{n}(K)$ by Lemma 8.9, and the dimension of the centralizer of $g$ in $\mathrm{Sp}_{n}(K)$ is $n/2$, and moreover $|C_{G}(g)|<4(q+1)^{n/2}$ by Lemma 7.11(iv). For $n\geq 8$ we see that $|M||C_{G}(g)|^{3}<|G|$ and the result follows. Assume then $n=4$. We have $q\geq 5$ in view of Lemma 8.1. We have $|g|=5$, and if $p\neq 5$ then $|C_{G}(g)|\leq q^{2}+1$, and if $p=5$ then $|C_{G}(g)|=2q^{2}$. The number of elements of $E.\mathrm{O}^{-}_{4}(2)=E.S_{5}$ of order 55 is $16\cdot 24=384$, so
@@
The other cases are similar. Assume $G=\mathrm{SU}_{n}(q)$ and $M$ of type $E.\mathrm{Sp}_{2m}(r)$ where $E$ is extraspecial of order $r^{2m+1}$ or $E=C_{4}\circ R$ where $R$ is extraspecial of order $2^{2m+1}$. Exactly as in the linear case, if $n=3$ (resp. $n=4$, resp. $n\geq 5$) then we choose $g$ of order 44 (resp. 55, resp. $r^{m}+1$), so $g$ is regular in $\mathrm{GL}_{n}(K)$. Finally, assume $G=\Omega^{+}_{n}(q)$ and $E.\Omega^{+}_{2m}(2)\leq M\leq E.\mathrm{O}^{+}_{2m}(2)$, where $n=2^{m}$ and $E=D_{8}\circ\cdots\circ D_{8}$ is an extraspecial group of order $2^{2m+1}$. Let $U:=E/Z(E)$ and let $g\in M$ be of order $2^{m}-1$, such that as an $\mathbb{F}_{2}\langle g\rangle$-module $U=A/Z(E)\oplus B/Z(E)$, where $A/Z(E)$ and $B/Z(E)$ are totally singular and irreducible. By Lemma 8.9, the dimension of the centralizer of $g$ in $\mathrm{GL}_{n}(K)$ is $n+2$. By Lemma 7.10(ii), it follows that the dimension of the centralizer of $g$ in $\mathrm{O}_{n}(K)$ is at most $(n+2)/2$, and we conclude with a similar calculation as above. ∎